# FIT-GNN: Faster Inference Time for GNNs Using Coarsening

### v2

@@ 5.2 Result and Analysis
Dataset Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
ZINC (Subset) $0.688$ $\mathbf{0.578}$ $0.657$
QM9 ($\mu$) $0.869$ $\mathbf{0.841}$ $0.948$
QM9 ($\Delta\epsilon$) $1.012$ $\mathbf{0.875}$ $0.969$
QM9 ($\mathrm{ZPVE}$) $1.091$ $\mathbf{0.818}$ $0.840$
QM9 ($U^{\mathrm{ATOM}}$) $1.078$ $0.754$ $\mathbf{0.630}$
Table 6: Graph regression tasks results on ZINC (Subset) and QM9. The metrics shown are in terms of normalized MAE (lower is better) with Gs-train-to-Gs-infer experimental setup. The Cluster Nodes method is used to append additional nodes to subgraphs.
Gc-train-to
-Gs-infer
Gs-train-to
-Gs-infer
Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
AIDS $0.792$ $0.810$ $0.836$ $0.812$ $\mathbf{0.843}$
PROTEINS $0.623$ $0.636$ $0.681$ $0.812$ $\mathbf{0.821}$
Table 7: Results for the graph classification task on AIDS and PROTEINS are presented. The metrics reflect accuracy; higher values indicate better performance. The table also compares two experimental setups.
Cora Chameleon
variation_neighborhoods $0.821\pm 0.003$ $0.803\pm 0.005$ $\mathbf{0.465\pm 0.005}$ $\mathbf{0.481\pm 0.006}$
kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
Table 3 shows the node regression error and standard deviation on coarsening ratios $0.1$ and $0.3$. All the experiments for node regression are shown with Cluster node method. With a higher coarsening ratio, more subgraphs are created. As the coarsening ratio increases, the size of each subgraph also reduces, and important neighboring node information gets lost. Hence, the performance decreases with an increase in the coarsening ratio.
Table 4 shows the results for node classification accuracy and standard deviation on different coarsening ratios on the Cora dataset with Cluster Nodes approach. With coarsening ratios $0.1$ and $0.3$, the observed performance is better than other coarsening ratios, and the inference time taken is also less. This motivates us to experiment on other datasets with lower coarsening ratios. Table 5 shows the results for node classification on different datasets using coarsening ratios $0.1$ and $0.3$. From the results, it can be observed that the accuracy is comparable to the baseline. Since experimenting on the full OGBN-Products dataset is not feasible due to memory constraints, we create a proxy graph by taking the maximum-sized communities with the number of nodes that sum up to $165000$. For the subset of the OGBN-Products graph, the performance is better with the baseline even with $0.56\%$ training nodes. For larger datasets like Pubmed and Physics Coauthor, a lower coarsening ratio doesn’t perform better than the baseline.
Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset. All the results show that the FIT-GNN model’s performance is comparable to the baselines. It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is high. It implies that finer subgraphs lose out global information about the molecule, which is necessary for the prediction.
Table 7 displays graph classification results using two experimental setups across two datasets. Training on coarsened graphs outperforms the baseline but discards important information. Conversely, training on subgraphs significantly improves predictions compared to the baseline.
PROTEINS ZINC (subset)
variation_neighborhoods $0.652$ $\mathbf{0.739}$ $\mathbf{0.629}$ $0.859$
algebraic_JC $\mathbf{0.783}$ $0.478$ $0.823$ $0.810$
kron $0.522$ $0.652$ $0.663$ $1.201$
heavy_edge $0.565$ $0.609$ $2.114$ $1.227$
variation_edges $0.652$ $0.565$ $1.548$ $1.883$
variation_cliques $0.696$ $0.652$ $0.689$ $\mathbf{0.742}$
Table 9: Results comparing various coarsening methods on PROTEINS and ZINC (subset) datasets. The metric for the PROTEINS dataset is accuracy (higher the better), and for the ZINC (subset) is normalized MAE (lower the better)
Table 8 and Table 9 show the comparison of performance with different coarsening algorithms used by Huang et al. (2021). We find that variation_neighborhoods is more consistent among all other coarsening algorithms, therefore we use it to report other results also.
5.3 Inference Time and Memory Comparison
FIT-GNN
Dataset Classical $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
@@ 5.3 Inference Time and Memory Comparison
Table 10: Inference time (sec) comparison for a single node using standard GNN as a baseline and FIT-GNN with two different coarsening ratios. Lower is better. Here, we used Cluster Nodes.
Figure 3: Comparison of inference time memory consumption (in MB) for Coauthor Physics Dataset between baseline and FIT-GNN
@@
Dastaset Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
@@
Table 11: Inference time (sec) comparison for the graph-level task. We use the Cluster Nodes method to add additional nodes to each subgraph and Gc-train-to-Gc-infer experiment setup
As mentioned in Section 4.3, our inference time is lesser than the standard GNN if certain conditions are met. Therefore, we empirically show reduction of time and space by our FIT-GNN model during inference. For recording the inference time, we use the Python time package to calculate the difference in time before and after the inference step.
In Table 10, we show the average time to predict for $1000$ nodes in the baseline and FIT-GNN models for node regression and classification tasks. The baseline model processes the entire graph, increasing inference time, especially for large graphs. In contrast, the FIT-GNN model only requires the relevant subgraph, resulting in faster predictions. For larger datasets like the OGBN-Products, we see up to $\mathbf{100\times}$ reduction in inference time. Table 17 in the Appendix shows the detailed comparison of memory consumptions for different datasets for both Extra Nodes and Cluster Nodes method along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{70\times}$ less memory than the baseline.
Table 11 compares the inference time of graph classification and graph regression task. We randomly select $1000$ graphs from the test split and infer them for baseline and FIT-GNN. We also append additional nodes to each subgraph using Cluster Nodes method. The table shows how our method is comparable and sometimes faster than the baseline. Also, we observe that the inference time increases with a higher coarsening ratio. Because, with a higher coarsening ratio, the number of nodes in $G^{\prime}$ increases, resulting in more edges. In table 16 in the Appendix, we show the estimated $\phi_{\textup{max}}$ value from the feature dimension of the nodes and the coarsening ratio. Combinations with negative numbers indicate the FIT-GNN-based approach won’t reduce the inference time.
Overall, the inference time and memory for all the tasks mentioned are drastically less than the classical approach while maintaining the performance mentioned in the previous section.
@@ 6 Conclusion
In this paper, we have focused on inference time and memory and presented a new way to utilize existing graph coarsening algorithms for GNNs. We have provided theoretical insight corresponding to the number of nodes in the graph for which the FIT-GNN model reduces the time and space complexity. Empirically, we have shown that our method is comparable to the uncoarsened baseline while being orders of magnitude faster in terms of the inference time and consuming a fraction of the memory. A few possible future directions involve studying directed and weighted graphs, focusing on the theoretical connections between Extra Nodes and Cluster Nodes.
Appendix A Dataset Description
@@
PROTEINS $1113$ $19$ $72$ $3$ $2$
AIDS $2000$ $7$ $16$ $38$ $2$
Table 12: Summary of datasets used for Graph classification Dataset
@@
QM9 $130831$ $8$ $18$ $11$ $19$
ZINC (subset) $10000$ $11$ $25$ $1$ $1$
Table 13: Summary of datasets used for Graph regression Dataset
@@
Cora $2708$ $5278$ $1433$ $7$
Citeseer $3327$ $4552$ $3703$ $6$
Pubmed $19717$ $44324$ $500$ $3$
DBLP $17716$ $52867$ $1639$ $4$
@@
$34493$ $247962$ $8415$ $5$
OGBN-Products $165000$ $4340428$ $100$ $47$
Table 14: Summary of datasets used for Node classification Dataset
@@
Chameleon $2277$ $31396$ $128$ $1$
Squirrel $5201$ $198423$ $128$ $1$
Crocodile $11631$ $170845$ $128$ $1$
Table 15: Summary of datasets used for Node regression Dataset
Dataset $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$
PROTEINS $6.999$ $0.333$ $-1.000$ $-1.571$
AIDS $322.000$ $82.000$ $34.000$ $13.429$
QM9 $79.000$ $19.000$ $7.000$ $1.857$
ZINC (subset) $-11.000$ $-4.333$ $-3.000$ $-2.429$
Table 16: Calculated $\phi_{\textup{max}}$ values for different datasets and coarsening ratios as per Lemma 1.
Appendix B More on Extra Nodes and Cluster Nodes
B.1 Extra Nodes
There are $3$ sets of nodes in $G_{i}$.
  • Nodes with 1-hop and 2-hop neighbours in $G_{i}$. Let us call this set $S_{1}$
$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-\mathcal{E}_{G_{i}}\right|$
The above entity will depend on the density of the subgraphs formed and the number of connections each subgraph shares with each other. An algorithm with an objective to reduce this entity for all subgraphs will lose the least amount of information when Extra Node method is used.
B.2 Cluster Nodes
Previously, according to Lemma 1, there is no information loss when using the $1$ layer of GNN and Extra Nodes method. It was also easy to quantify in terms of the number of nodes. However, it is different for Cluster Nodes. Let us discuss the issues first.
$\bullet$ Only a weighted version of node information is shared with the subgraph. Suppose $v_{c},v_{d}\in G_{i}$ is connected to $v_{a},v_{b}\in G_{j}$. Then the information contributed by these nodes is $\frac{d(v_{a})x_{a}+d(v_{b})x_{b}}{\sum_{p}d(v_{p})}$. Here $d(v_{p})$ represents the degree of node $v_{p}$, and $x_{p}$ represents the feature of node $v_{p}$.
$\bullet$ Other node information will also be shared which is $\frac{\sum_{p\neq v_{a},v_{b}}d(v_{p})x_{p}}{\sum_{p}d(v_{p})}$. This will capture further dependencies.
Appendix C Experiment Details (Parameters and Device Configuration)
@@
For node classification and node regression tasks, we use Adam Optimizer with a learning rate of $0.01$ and $L_{2}$ regularization with $0.0005$ weight. We use Adam Optimizer with a learning rate of $0.0001$ and $L_{2}$ regularization with $0.0005$ weight for graph level tasks. For both coarsened graph and set of subgraph based model, we set epochs to 300 and the number of layers of GCN to 2 for training. We set the hidden dimensions for each layer of GCN to $512$.
@@
The device configurations are Intel(R) Xeon(R) Gold 5120 CPU @ $2.20$GHz, $256$GB RAM, NVIDIA A100 $40$GB GPU. We use Pytorch Geometric to to train our models.
Appendix D Results
@@
This section shows the memory consumption for different datasets with different methods. Also, we show how the FIT-GNN model performs with different methods of appending additional nodes to subgraphs at different coarsening ratios.
FIT-GNN
Dataset Split $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$ Baseline
Cora cluster $1.310$ $0.648$ $0.694$ $0.868$ $15.743$
extra $2.072$ $1.134$ $1.017$ $1.017$
Citeseer cluster $31.559$ $2.148$ $1.253$ $1.686$ $49.488$
extra $31.559$ $3.114$ $1.788$ $1.788$
Pubmed cluster $0.613$ $0.569$ $0.571$ $0.573$ $41.227$
extra $0.798$ $0.590$ $0.590$ $0.590$
DBLP cluster $6.067$ $2.940$ $1.847$ $1.966$ $118.121$
extra $13.907$ $4.663$ $2.649$ $2.274$
Physics Coauthor cluster $24.025$ $13.088$ $13.388$ $14.353$ $1169.521$
extra $63.146$ $22.434$ $22.434$ $22.434$
Chameleon cluster $0.212$ $0.246$ $0.291$ $0.438$ $2.198$
extra $0.653$ $0.653$ $0.653$ $0.592$
Crocodile cluster $1.183$ $1.188$ $1.735$ $1.866$ $11.562$
extra $4.368$ $4.368$ $4.368$ $3.134$
Squirrel cluster $1.507$ $1.607$ $1.903$ $3.009$ $9.075$
extra $7.275$ $7.275$ $7.275$ $6.953$
OGBN-Products cluster $6.047$ $5.613$ $8.636$ $208.029$

### v3

@@ 6.2 Inference Time and Memory
Dataset Baselines $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
@@
(a) Node-Level Tasks
@@
Dastaset Baselines $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
@@
(b) Graph-Level Tasks
Table 8: Inference time (sec) (Lower is better) comparison for Node-Level tasks and Graph-Level tasks. We show single-node prediction time using two different coarsening ratios for Node-Level tasks. Here, we used Cluster Nodes for the node-level task. We used Gc-train-to-Gc-infer experimental setup for the graph-level tasks.
[…]
Figure 3: GPU memory consumption (log scale) for FIT-GNN (Cluster Node) at different reduction ratios $r$ and Baseline, during inference.
In Table 8(a), we show the average time to predict for $1000$ nodes in the baseline and FIT-GNN models for node regression and classification tasks. The baseline model processes the entire graph, increasing inference time, especially for large graphs. In contrast, the FIT-GNN model only requires the relevant subgraph, resulting in faster predictions. For larger datasets like the subset of OGBN-Products, baseline inference is not feasible with our computational capacity. Therefore, to show a comparison, we take a subset of OGBN-Products with $165000$ nodes and $4340428$ edges. We see up to $\mathbf{100\times}$. Section C elaborates more on the inference time for datasets with a larger number of nodes. Traditional methods fail to infer for very large datasets due to memory constraints. However, our approach can give an inference on the whole graph. Table 11 in the Appendix shows the detailed comparison of memory consumption for different datasets for both Extra Nodes and Cluster Nodes methods, along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{100\times}$ less memory than the baseline.
Table 8(b) compares the inference time of graph classification and graph regression tasks. We predict for randomly selected $1000$ graphs from the test split. The table shows how our method is comparable and sometimes faster than the baseline. Also, we observe that the inference time increases with a higher coarsening ratio. Because, with a higher coarsening ratio, the number of nodes in $G^{\prime}$ increases, resulting in more edges. Overall, the inference time and memory for all the tasks mentioned are drastically less than the baselines while maintaining the performance.
@@ 7 Conclusion
In this paper, we have focused on inference time and memory and presented a new way to utilize existing graph coarsening algorithms for GNNs. We have provided theoretical insight corresponding to the number of nodes in the graph for which the FIT-GNN model reduces the time and space complexity. Empirically, we have shown that our method is comparable to the uncoarsened baseline while being orders of magnitude faster in terms of inference time and consuming a fraction of the memory. A few possible future directions involve studying directed and weighted graphs, focusing on the theoretical connections between Extra Nodes and Cluster Nodes, and tackling real-world data-intensive challenges, for example, weather forecasting.
Appendix
Appendix A More on Extra Nodes and Cluster Nodes
A.1 Extra Nodes
Lemma 4.1. Models with $1$ layer of GNN cannot distinguish between $G$ and $\mathcal{G}_{s}$ when Extra Nodes method is used.
Proof.
There are $2$ sets of nodes in $G_{i}$ as follows:
  • The set $S_{1}$ of nodes with 1-hop neighbours in $G_{i}$.
  • The set $S_{2}$ of nodes with not all 1-hop neighbors in $G_{i}$
[…]
$\Rightarrow\left|\mathcal{E}_{G_{i}}\right|=\mathcal{I}^{1}_{i}$
Hence, when Extra Nodes is used, 1 Layer GNN model cannot distinguish $G$ and $\mathcal{G}_{s}$. ∎
To understand the information loss when taking $2$ Layers of GNN, we divide $G_{i}$ into $3$ sets of nodes.
  • Nodes with 1-hop and 2-hop neighbours in $G_{i}$. Let us call this set $S_{1}$
$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-V(G_{i})-\mathcal{E}_{G_{i}}\right|$
The above entity will depend on the density of the subgraphs formed and the number of connections each subgraph shares with each other. An algorithm with an objective to reduce this entity for all subgraphs will lose the least amount of information when Extra Nodes method is used.
A.2 Cluster Nodes
Previously, according to Lemma 4.1, there is no information loss when using the $1$ layer of GNN and Extra Nodes method. It was also easy to quantify in terms of the number of nodes. However, it is different for Cluster Nodes. Let us discuss the issues first.
  • Only a weighted version of node information is shared with the subgraph. Suppose $v_{c},v_{d}\in G_{i}$ is connected to $v_{a},v_{b}\in G_{j}$. Then the information contributed by these nodes is $\frac{d_{a}x_{a}+d_{b}x_{b}}{\sum_{p}d_{p}}$. Here $d_{p}$ represents the degree of node $v_{p}$, and $x_{p}$ represents the feature of node $v_{p}$.
  • Other node information will also be shared which is $\frac{\sum_{p\neq v_{a},v_{b}}d_{p}x_{p}}{\sum_{p}d_{p}}$. This will capture further dependencies.
Appendix B More Algorithms
Algorithm 3 Training GNN (Train on $G^{\prime}$) [8]
1:$G=(V,E,X)$; Labels $Y$; Node-Model $M$; Loss $\ell$; Number of Layers $L$
2:Apply a coarsening algorithm on $G$, and output a normalized partition matrix $P$;
3:Construct Coarsened graph $G^{\prime}$ using $P$;
4:Construct feature matrix for $G^{\prime}$ by $X^{\prime}=P^{T}X$;
5:Construct labels for $G^{\prime}$ by $Y^{\prime}=\arg\max(P^{T}Y)$ ;
[…]
8:Train $M$ to minimize LossLoss;
Algorithm 4 Node-Model(L,A,D,X)(L,A,D,X)
1:Number of Layers $L$; AA; $D$; $X^{(0)}=X$;
2:for $i=1$ to $L$ do
3:  $X^{(i)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(i-1)}\mathcal{W}^{(i-1)})$ $\quad\#$ Equation 1
4:end for
[…]
6:return $Z$
1:Number of Layers $L;A^{\prime};D^{\prime};X^{(0)}=X^{\prime}$
2:for $i=1$ to $L$ do
[…]
4:end for
[…]
7:return $Z$
Appendix C More on Time and Space complexity
Figure 4: The plot shows the feasibility region (green) corresponding to the coarsening ratio and Number of nodes (In order of Millions) in the graph for a fixed feature dimension of $d=200$.
In Lemma 4.2, we provide the conditions that need to be satisfied for better space and time complexities. Figure 4 shows that those conditions can be satisfied for any dataset with nodes in order of millions.
For node-level tasks, we show that the inference phase will take less time for single node prediction. To show that the time difference between the time taken by the classical approach and ours will increase as the number of nodes increases, we can formulate the time difference as follows:
$\mathcal{T}(n)=\underset{\textup{Baseline Complexity}}{\underbrace{n^{2}d+nd^{2}}}-\underset{\textup{ FIT-GNN Complexity}}{\underbrace{\max_{i}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]}}$ (3)
In the above equation, we use the notations already defined in Section 4.3. We can also assume that $\max_{i}^{k}n_{i}=\frac{n}{\alpha}$, where $\alpha>1$ is some constant. Then we can rewrite the Equation 3 as:
$\mathcal{T}(n)=n^{2}d+nd^{2}-(\frac{n}{\alpha}+\phi_{\max})^{2}d+(\frac{n}{\alpha}+\phi_{\max})d^{2}$
We can observe that the above equation is strictly increasing for $n>0$. Hence the time difference also increases as the number of nodes in a dataset increases.
Table 9: Upper Bound of $\phi_{\textup{max}}$ values for different datasets and coarsening ratios.
Dataset $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$
PROTEINS $6.999$ $0.333$ 1.000-1.000 1.571-1.571
AIDS $322.000$ $82.000$ $34.000$ $13.429$
QM9 $79.000$ $19.000$ $7.000$ $1.857$
ZINC (subset) 11.000-11.000 4.333-4.333 3.000-3.000 2.429-2.429
Appendix D Dataset Description
@@
PROTEINS $1113$ 1919 $72$ 33 22
AIDS $2000$ 77 $16$ $38$ 22
(a) Summary of datasets used for Graph classification Dataset
@@
QM9 $130831$ 88 $18$ $11$ 1919
ZINC (subset) $10000$ $11$ $25$ 11 11
(b) Summary of datasets used for Graph regression Dataset
@@
Cora $2708$ $5278$ $1433$ 77
Citeseer $3327$ $4552$ $3703$ 66
Pubmed $19717$ $44324$ $500$ 33
DBLP $17716$ $52867$ $1639$ 44
@@
$34493$ $247962$ $8415$ 55
OGBN-Products $2449029$ $61859140$ $100$ $47$
(c) Summary of datasets used for Node classification Dataset
@@
Chameleon $2277$ $31396$ $128$ 11
Squirrel $5201$ $198423$ $128$ 11
Crocodile $11631$ $170845$ $128$ 11
(d) Summary of datasets used for Node regression Dataset
Appendix E Experiment Details (Parameters and Device Configuration)
@@
For node classification and node regression tasks, we use Adam Optimizer with a learning rate of $0.01$ and $L_{2}$ regularization with $0.0005$ weight. We use Adam Optimizer with a learning rate of $0.0001$ and $L_{2}$ regularization with $0.0005$ weight for graph level tasks. For both coarsened graph (Gc-train) and set of subgraph based (Gs-train) model, we set epochs to 300 and the number of layers of GCN to 2 for training. We set the hidden dimensions for each layer of GCN to $512$.
@@
The device configurations are Intel(R) Xeon(R) Gold 5120 CPU @ 2.202.20GHz, 256256GB RAM, NVIDIA A100 4040GB GPU. We use Pytorch Geometric to to train our models.
Appendix F Results
@@
Table 11 shows the maximum GPU memory consumption during inference for different datasets with different coarsening ratios and methods of appending nodes. Also, in Table 12 we show how the FIT-GNN model performs on Node Regression datasets for different coarsening ratios with Cluster Nodes method under Gs-train-to-Gs-infer setup.
Dataset Appending Nodes FIT-GNN Baseline
$\mathbf{=0.1} r$$\mathbf{=0.3}$ $\mathbf{=0.5} r$$\mathbf{=0.7}$
Chameleon Cluster Node $0.201$ $0.235$ $0.277$ $0.418$ $2.078$
Extra Node $0.622$ $0.622$ $0.622$ $0.564$
Crocodile Cluster Node $1.127$ $1.132$ $1.654$ $1.779$ $10.937$
Extra Node $4.586$ $4.586$ $4.586$ $4.586$
Squirrel Cluster Node $1.436$ $1.532$ $1.815$ $2.869$ $8.614$
Extra Node $7.190$ $7.190$ $7.190$ $7.190$
Cora Cluster Node $1.249$ $0.618$ $0.661$ $0.827$ $14.992$
Extra Node $1.906$ $1.103$ $0.970$ $0.970$
Citeseer Cluster Node $30.097$ $2.048$ $1.195$ $1.607$ $47.170$
Extra Node $30.097$ $2.490$ $1.835$ $1.835$
Pubmed Cluster Node $0.584$ $0.542$ $0.544$ $0.546$ $39.166$
Extra Node $0.778$ $0.562$ $0.562$ $0.562$
DBLP Cluster Node $5.785$ $2.803$ $1.761$ $1.875$ $112.514$
Extra Node $28.271$ $6.008$ $2.308$ $2.207$
Physics Coauthor Cluster Node $22.911$ $12.481$ $12.767$ $13.688$ $1115.079$
Extra Node $41.066$ $19.936$ $19.936$ $19.936$
OGBN-Products Cluster Node $28.399$ $2840.706$